OpenStudy (anonymous):

Prove the identity: tan2x = (2tan x)/(1-tan^2 x)

OpenStudy (alfie):

Alright, I just worked this out: \[\frac{sen2x}{\cos2x} \rightarrow \frac{2senxcosx}{\cos^2x-sen^2x}\] We are free to multiply both denominator and denominator for the same quantity. \[\frac{\frac{cosx2senxcosx}{cosx}}{\frac{\cos^2x (\cos^2x - sen^2x)}{\cos^2x}}\] Now, doing few calcs you get: \[\frac{2tanx}{\frac{\cos^2x}{\cos^2x} - \frac{sen^2x}{\cos^2x}}\] Which eventually goes tO: \[\frac{2tanx}{1-\tan^2x}\] :) Bye.